What is the concentration of HI in a solution with pH=4.86?

Answer choice:

a. 1.38*10^-5M

b. 2.69*10^-5M

c. 7.25*10^-10

d. 3.72*10^-10
Answer #1


we have below equation to be used:
pH = -log [H+]
4.86 = -log [H+]
log [H+] = -4.86
[H+] = 10^(-4.86)
[H+] = 1.38*10^-5 M

HI is strong acid
So,
[HI] = [H+] = 1.38*10^-5 M

Answer: 1.38*10^-5 M